# Audio xLSTMs: Learning Self-supervised audio representations with xLSTMs

### v1

@@ III-B Implementation details
Pretraining: All the proposed models are pretrained on randomly cropped 2-second audio clips, which corresponds to a [200, 80]-dim (time and frequency, resp.) input. Our default configuration consists of $l=12$ number of stacked mLSTM blocks with a model feature dimension $d_{m}=192$, same as those of ViT-Tiny [23] encoder, but we also evaluate Small ($d_{m}=$384, $l$=12) and Base ($d_{m}$=768, $l$=12) encoder configurations. The default expansion factor $E_{f}=2$ as specified in [20, 21] is used in the underlying mLSTM layers. By default, a patch embedding layer that computes ($4\times 16$) shaped non-overlapping patches is used in all our experiments. All models are trained for 100 epochs with a batch size of 1024 and a weight decay of 0.05. AdamW optimizer with a linear warmup for 10 epochs followed by a cosine learning rate decay schedule is used. No data augmentations were used. We follow hyperparameter recommendations from previous work [19, 7].
@@
Downstream evaluation: Following the HEAR protocol, we extract fixed sized feature vectors independent of the input audio duration by taking the mean over time across 2-second audio chunks. We then train a single hidden layer MLP classifier with 1024 neurons for each task, using the official hear-eval-kit accompanying the HEAR benchmark. All experiments are repeated with 10 different random seeds.
@@
Aggregated Performance Metric: We use the aggregated normalized score as proposed by [27] to compare evaluated approaches across the proposed list of downstream tasks. For a model $m$, overall score $s(m)\in[0.,100.]$ is given as: $s(m)=\frac{1}{|T|}\sum_{t\in T}\frac{x_{t}(m)-min_{t}}{max_{t}-min_{t}}*100$, where $x_{t}(m)$ denotes performance of the model $m$ on task $t$, and $min_{t}$ and $max_{t}$ represent the worst and the best performance across all models on the task, thus taking into account the relative performance amongst all evaluated representations.
IV Results
IV-A Comparison with existing works
@@
Table I shows how the proposed AxLSTM models fare against recent audio representations. Directly comparable SSAST and AxLSTM feature representations are colour coded and have identical feature embedding sizes, however, feature vector sizes extracted using other referred methods can vary. While sub-optimal, it is infeasible to retrain all pretrained representations to have the same embedding sizes, and our evaluation protocol is in line with recent frameworks for evaluating self-supervised audio representations [24, 28]. SSAST [7] represents the official SSAST released model, which was pretrained on AudioSet and LibriSpeech datasets with the masked prediction + reconstruction multitask objective. We can see that AxLSTM models consistently outperform their transformer based SSAST counterparts by a considerable margin, while having over 45% fewer parameters, with AxLSTM-Base configuration yielding a 30% relative improvement in aggregate performance (83.1±0.2 v/s 69.2±0.3). While Mamba based SSAM [19] as well as Masked Autoencoder based approaches [27] perform better, it is worth noting that the proposed AxLSTM models have about 35% and 50% fewer parameters, respectively. Overall, we can conclude that AxLSTMs performs very favourably compared to popular audio representations.
IV-B Ablations
Expansion Factor: We investigate how AxLSTM models scale with different expansion factors ($E_{f}$). Table II shows that while using $E_{f}=3$ improves performance, $E_{f}=4$ leads to worse performance, suggesting overfitting.
TABLE II: Performance impact of expansion factor ($E_{F}$)
@@ IV-B Ablations
AxLSTM-Tiny 2 2.9 M 67.0±0.3
AxLSTM-Tiny 3 4.3 M 68.9±0.4
AxLSTM-Tiny 4 5.8 M 63.4±0.2
Patch Size: To investigate how AxLSTM models perform compared to SSAST models with changing number of input patches, we pretrain AxLSTM-Tiny models with 3 patch sizes: $(4,8)$, $(4,16)$, and $(8,16)$. AxLSTM scales better with increasing number of patches, with smaller patch sizes leading to better performance (Table III). We expect expansion factor $E_{f}=3$ paired with a patch size of $(4,8)$ to further improve performance.
TABLE III: Patch size ablations with the Tiny configuration
@@
SSAST-Tiny $(8,16)$ 125 47.3±0.3
AxLSTM-Tiny $(8,16)$ 125 55.3±0.3
SSAST-Tiny $(4,16)$ 250 56.0±0.2
AxLSTM-Tiny $(4,16)$ 250 67.0±0.3
SSAST-Tiny $(4,8)$ 500 54.1±0.4
AxLSTM-Tiny $(4,8)$ 500 68.3±0.3
Other design choices: There are several design choices that impact the modeling performance of AxLSTMs. In Table IV, we evaluate the impact of two of the most important of these factors using the AxLSTM-Tiny configuration. We observe that the absence of exponential gating for the input and forget gates in mLSTM ends up improving performance. This can be attributed to the fact that for the given input sequence length (250 tokens), revising storage decisions is not a bottleneck. Further, we observe that removing sequence flipping in even numbered blocks [21] significantly boosts performance, highlighting that bidirectional processing of patches is not as important in the given framework as it is for vision, which is in line with observations made in previous works [19].
TABLE IV: Evaluating other design choices for AxLSTM-Tiny
$s(m)$
default 67.0±0.3
default, no exponential gating 68.2±0.2
default, no sequence flipping 70.2±0.3
V Conclusion
This work presents extended long short-term memory models for audio (AxLSTMs), an approach for learning self-supervised general-purpose audio representations from masked audio spectrograms which extends the recently proposed xLSTM neural architecture. Using the AudioSet corpus for pretraining, we evaluate AxLSTMs against directly comparable Self-supervised Audio Spectrogram transformer (SSAST) baselines on ten varied downstream audio recognition tasks. Our empirical analysis shows that AxLSTMs offer up to 20% improvement in relative performance over their transformer based counterparts while having up to 45% fewer parameters, thus establishing the viability of xLSTM approaches for learning self-supervised audio representations.
Acknowledgment
We would like to thank Prof. Lars Kai Hansen, Technical University of Denmark, for valuable feedback. We are also grateful to the TPU Research Cloud Program, a Google Research Initiative, for providing TPU devices used in this project, as well as the Danish e-Infrastructure Consortium (DeiC) for enabling access to the LUMI supercomputer.

### v3

@@ 3.2 Implementation details
Pretraining: All the proposed models are pretrained on randomly cropped 2-second audio clips, resulting in an input spectrogram of shape $[\mathtt{200}\times\mathtt{80}]$. Our default configuration consists of $l=12$ number of stacked mLSTM blocks with a model feature dimension of $d_{m}=192$, the same as those of a ViT-Tiny [20] encoder, but we also evaluate Small ($d_{m}=$384, $l$=12) and Base ($d_{m}$=768, $l$=12) encoder configurations. By default, a patch embedding layer that computes $[4\times 16]$ shaped non-overlapping patches (along time and frequency, respectively) is used in all our experiments. While [16, 17] suggest a default expansion factor of $E_{f}=2$, in early ablations we found that $E_{f}=3$ performed the best (Section 4.2). Ablation experiments also revealed that disabling the flip operation performed better than enabling flip in alternating blocks. We build on hyperparameter recommendations from previous work [6, 14]: all models are trained for 100 epochs with a batch size of 1024 and a weight decay of 0.05. AdamW optimizer with a linear warmup for 10 epochs followed by a cosine learning rate decay schedule is used. No data augmentations were used.
@@
Downstream evaluation: Following the HEAR protocol, we extract fixed-sized feature vectors independent of the input audio duration by taking the mean over time across 2-second audio chunks. We then train a single hidden layer MLP classifier with 1024 neurons for each task, using the official hear-eval-kit accompanying the HEAR benchmark. All experiments are repeated with 10 different random seeds, and 95% confidence intervals are reported. It is worth noting that due to the number of experiments conducted, we use a restricted hyperparameter grid (same as [14, 22]) compared to the hear-eval-kit.
@@
Aggregated Performance Metric: We use the aggregated normalized score as proposed by [22] to compare evaluated approaches across the proposed list of downstream tasks. For a model $m$, overall score $s(m)\in[0,100]$ is given as: $s(m)=\frac{1}{|T|}\sum_{t\in T}\frac{x_{t}(m)-min_{t}}{max_{t}-min_{t}}*100$, where $x_{t}(m)$ denotes performance of the model $m$ on task $t$, and $min_{t}$ and $max_{t}$ represent the worst and the best performance across all models on the task, thus taking into account the relative performance amongst all evaluated representations.
4 Results
4.1 Comparison with existing works
@@
Table 2 shows how the proposed AxLSTM models fare against recent audio representations. Directly comparable SSAST and AxLSTM feature representations have identical feature embedding sizes, however, feature vector sizes extracted using other referred methods can vary. While sub-optimal, it is infeasible to retrain all pretrained representations to have the same embedding sizes, and our evaluation protocol is in line with recent frameworks for evaluating self-supervised audio representations [23, 34]. SSAST [6] represents the officially released model trained on AudioSet+LibriSpeech, whereas underlined SSAST representations were pretrained by us. We can see that AxLSTM models consistently outperform their transformer-based SSAST counterparts by a considerable margin, while having over 23% fewer parameters, with AxLSTM-B yielding over 25% relative improvement in aggregate performance (86.6±0.2 v/s 68.7±0.3). Despite being based on the older SSAST framework, our AxLSTM-B model matches the more recent masked modeling-based approaches such as AudioMAE [7], BEATS-iter3 [19] and MSM-MAE-208 [21] in overall performance, while having over 25% fewer parameters. The recent Mamba-based SSAM [14] models outperform comparable AxLSTM models in overall performance, however, it’s worth noting that AxLSTM models outperform SSAM models in music and pitch perception tasks (BO, Mri-T, NSynth-5h) whereas SSAM models shine in Speech-based and Audio classification tasks, performing notably better for emotion recognition (CD) and spoken language identification (VL). Finally, the recent MWMAE-Base [22] model with local-global self-attention outperforms every evaluated approach, including AxLSTM-B, albeit with notably more parameters. Overall, while the AxLSTM model doesn’t establish a new state-of-the-art, it significantly outperforms SSAST-based transformer baselines while matching the performance of several other newer masked modeling approaches while possessing fewer parameters, thus demonstrating the viability of xLSTMs to learn self-supervised audio representations.
4.2 Ablations
This section covers ablations for investigating the impact of key configurable hyperparameters. Unless stated otherwise, experiments use the Tiny configuration with an expansion factor $E_{f}=2$ and flipping enabled in alternating mLSTM blocks.
Expansion Factor: This hyperparameter affects the size and modeling capacity of mLSTM blocks. Table 3 shows that $E_{f}=3$ offers better overall performance than the default $E_{f}=2$ (suggested in [16, 17]). However, $E_{f}=2$ still outperforms the SSAST baseline, offering 20% better performance with over 45% fewer parameters.
Table 3: Performance impact of expansion factor ($E_{F}$)
@@ 4.2 Ablations
SSAST-T - 5.4 M 55.6±0.2
AxLSTM-T 2 2.9 M 66.9±0.3
AxLSTM-T 3 4.3 M 68.9±0.4
AxLSTM-T 4 5.8 M 63.4±0.2
Patch Size: The patch size hyperparameter governs the sequence length as well as the time-frequency resolution of the patches that are processed by the underlying sequence modeling blocks. We pretrain and compare SSAST and AxLSTM Tiny configurations with 3 patch sizes: $(4,8)$, $(4,16)$, and $(8,16)$. From Table 4, we can see that AxLSTM scales better with increasing number of patches, with smaller patch sizes continually leading to better performance, whereas SSAST-T experiences a drop in performance at patch size of $(4,8)$, suggesting that AxLSTMs scale better with sequence length in terms of modeling performance compared to standard transformers.
Table 4: Patch size ablations with the Tiny configuration
@@
SSAST-T $(8,16)$ 125 47.0±0.3
AxLSTM-T $(8,16)$ 125 55.2±0.3
SSAST-T $(4,16)$ 250 55.6±0.2
AxLSTM-T $(4,16)$ 250 66.9±0.3
SSAST-T $(4,8)$ 500 53.8±0.3
AxLSTM-T $(4,8)$ 500 68.2±0.3
Bidirectional operation and exponential gating: In Table 5, we evaluate the impact of bidirectional modeling (controlled through the flip operation) and replacing exponential gating with a sigmoid gate. When done one at a time, disable both flip operation in alternating xLSTM blocks and the exponential gate improve performance compared to when both of these operations are enabled, with the former leading to a greater performance improvement. This suggests that bidirectional modeling is not as important for downstream audio classification as it is for vision [17], which is inline with observations made in previous work [14]. However, disabling both exponential gating and flipping together does not further improve performance, instead it performs worse than just disabling sequence flipping in alternative blocks. These results suggest that changes at the architectural level to accommodate bidirectional modeling and exponential gating might be warranted. It is also worth noting that the best performing AxLSTM-T model with $E_{f}=2$ outperforms the SSAST-T baseline by 26% but has 45% fewer parameters.
Table 5: Evaluating the impact of exponential gating and flipping on overall performance
Model $E_{f}$ Exp gate Flip $s(m)$
AxLSTM-T 2 66.9±0.3
AxLSTM-T 2 70.1±0.3
AxLSTM-T 2 68.2±0.2
AxLSTM-T 3 68.8±0.4
AxLSTM-T 3 70.7±0.2
AxLSTM-T 3 69.6±0.2
5 Conclusion
This work presents AxLSTMs, an approach that evaluates the viability of the recently proposed xLSTM model for learning self-supervised audio representations from masked audio spectrograms. Extensive empirical analysis shows that AxLSTMs significantly outperform comparable SSAST baselines on ten varied downstream audio recognition tasks. AxLSTM-B, which is our best performing model, offers over 25% better performance than SSAST-B while having up to 23% fewer parameters, and even matches the performance of several newer masked modeling approaches such as MSM-MAE-208 and BEATs. On the other hand, AxLSTM-T models with $E_{f}=2$ can offer over 25% relative improvement compared to an SSAST baseline, while having over 45% fewer parameters, which is a great performance/complexity tradeoff. Overall, our results demonstrate the viability of xLSTMs for learning self-supervised audio representations.